Previous

Calendar

August

2015

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

26

JD207

27

JD208

28

JD209

29

JD210

30

JD211

31

JD212

1

JD213

2

JD214

3

JD215

4

JD216
SpB Returns

5

JD217

6

JD218
Site Visit (SpB): Details , Photos (15)

Span cylinder at 750 psi. N2 at 400 psi

at 7:00 MST, rebooted LI-7200 (via network)

from 10:20-10:30 MST, added PAR-LITE sensor (PAR2.up) to top tower level

at 11:50 MST, started ops80: changed order of Ttcs and Accel1 on dsm chan 1,90 (started mark_06.csi)

at 14:00 MST, re-started the 2.5m LI-7500A

from 12:00-14:00 MST, replaced the knf prof micropump

from 15:00-15:30 MST, for Ttc.csat4; added fine-wire; moved to the end of boom (co-located with 15m CSAT)

swapped SD cards from towerCAM, cloudCAMwest and cloudCAMeast

MRS Shuttle

Windy and cool (mix of clouds/sun)

7

JD219

8

JD220

9

JD221
at approx 12:00 MST, the power supply for the LI-7200 flow module and LI-7500A died

10

JD222

11

JD223
at 4:00 MST, cloudCAMwest stops working

site visit by Mark/Ethan?

12

JD224

13

JD225
site visit by DB

14

JD226
Site Visit (SpB, DB, DE): Details , Photos (20)

Span cylinder at 700 psi. N2 at 300 psi

the power supply for the LI-7200 flow module and LI-7500A was dead (no flow in LI-7200!)

at 10:45 MST, started ops81: removed Accel1 on dsm chan 1,90 (started mark_07.csi)

from 11:15 MST, turned off prof knf micropump (need to replace Nupro filter)

from 12:00-12:15 MST, swapped thick with fine-wire for Ttc.csat1; also moved closer to CSAT path (co-located with 5m CSAT)

swapped SD cards from towerCAM

SpB Jeep

Calm, some big Cu clouds

site visit by DB

15

JD227

16

JD228

17

JD229

18

JD230

19

JD231

20

JD232

21

JD233
Site Visit (SpB, DB, H): Details , Photos (29)

Span cylinder at 650 psi. N2 at 100 psi --> 300 psi

at approx 8:30 MST, removed cc142673 (100 psi) re-started cc316422 (n2) at 300 psi

at 9:30 MST, replaced the battery charger for the the LI-7200/LI-7550/flow module and LI-7500A (OP2)

from 11:15-11:45 MST, restarted prof system (needed to adjust flow rate!!)

from xx:xx-xx:xx MST, repositioned the LI-7500 (moved closer to CSAT, and upside-down)

Just before leaving C1: noticed the prof logger data are not being recorded by the data system

in PM, checked the connection for Ttc.csat3 at the logger (seemed ok)

swapped SD cards from cloudCAMwest and cloudCAMeast (note: cloudCAMwest is DEAD!)

MRS Shuttle

Very hazy/smokey; some wind gusts in afternoon

22

JD234
Site Visit (DB): Details , Photos (6)

DB finished up sap flow sensors

23

JD235
Site Visit (SpB): Details , Photos (3)

Span cylinder at ? psi. N2 at 250 psi

at 11 MST, found that prof logger was unplugged! plugged it back in, then working again

SpB Wagon

Still hazy/smokey

24

JD236

25

JD237
during the night: something strange with LI-7200 CO2? Flow rate? Dirty optics (due to smoke)?

26

JD238

27

JD239
Site Visit (SpB, AR, DA): Details , Photos (30)

Span cylinder at 600 psi. N2 at 100 --> 2000 psi

at approx 12:30 MST, removed cc316422 (100 psi) started using cc153282 (n2) at 2000 psi

more laptop issues (plumwood was acting odd?)

updated firmware for LI-7500A (OP1) (from 7.0.1 to 7.5.0); now all three IRGAs are using 7.5.0

from 13:25-13:30 MST, swapped thermocouple for Ttc.csat3 (still has dropouts!)

in PM: shot boom angles for NCAR CSATs (using SO sextant)

in PM: AR/DA were installing FLIR camera

swapped SD card from towerCAMtree

CU prius

Overcast with light showers on/off

28

JD240
site visit by AR/DA

29

JD241
site visit by AR/DA (finish camera setup)

(Full "Green Corn" Moon)

30

JD242

31

JD243

1

JD244

2

JD245

3

JD246

4

JD247
Site Visit (SpB): Details , Photos (32)

Span cylinder at 550 psi. N2 at 1820 psi

at 12:00 MST, a slight decrease in sample flow rate of LI-6251 (prof)

at 12:30 MST, changed bandpass in 2.5m LI-7500 from 20-Hz to 5-Hz

from 13:00-13:02 MST, adjusted span/zero flow in LI-6262

at 14:00-14:30 MST, started ops82: added T-type thermocouple to mark logger (Ttc.Ttype, started mark_08.csi)

from 14:30-14:40 MST, in mark logger: swapped diff chan 9 (Ttc.csat3) with diff chan 8 (Ttc.csat2)

SpB Jeep

Overcast with light showers

5

JD248